1). The result of Java expression 5*7>3*(5+2) && 3*5-4<6 is ____
a. 35>35&&9<6
b. true && false
c. false
d. False

(2). You save text files containing Java language source code using the file extension___
a. .java
b. .class
c. .txt
d. .src
  
(3). Which of the following is not a primitive data type in the Java programming language?
a. boolean
b. byte
c. int
d. Point

(4). Given the Java statement int[][] vari = new int[5][7]; the value of vari[4].length is
a. 4
b. 5
c. 6
d. 7

(5). To use any of the prewritten classes besides those in the java.lang package, you must___
a. use the entire path with the class name
b. import the class
c. import the package of which the class you are using is a part
d. use any of these methods

(6). Assuming x = 25, y = -1, z = 32, the expression !( x / 2 == 12 ) || ( z*3 == 96 ); gives ___
a. False
b. true
c. false
d. True

(7). Given String str1 = “Java is useful!”; Statement System.out.println(str1.charAt(14)); displays ___
a. 14
b. !
c. l
d. error message
  
(8). Given a static method public static void swap(int x, int y){ int temp = x; x = y; y = temp; }, if x=3, y=5; then the value of x and y after the method call swap(x,y) is_____;
a. 3, 5
b. 5, 3
c. 3, 3
d. 5, 5

(9). A class header must contain___
a. an access modifier
b. the keyword class
c. the keyword static
d. the keyword public

(10). Which of the following expression refers the last element of array arr?
a. arr[last]
b. arr[arr.length]
c. arr[arr.length – 1]
d. arr[end]

(11). Given Java statement int[] arr = new int[10]; the last element is___
a. 10
b. arr[10]
c. arr[9]
d. arr[0]

(12). Given Java statement Random rand = new Random( ); which of the following Java statements will
generate a random integer between 2 and 5 is
a. rand.nextInt(4)+2;
b. rand.nextInt(5) +2;
c. rand.nextInt(6) +2;
d. rand.nextInt(2,5);

(13). You can declare two variables with the same name as long as___
a. they appear within the same block
b. they are assigned different values
c. they are of different types
d. their scopes do not overlap

(14). Given Java statement int[] arr = new int[10]; the statement System.out.println(arr) gives__
a. 10
b. 9
c. compiler error
d. The memory location where array arr is stored.

(15). If str is a String, ______ gives the total number of characters in the string str.
a. str.length
b. str.length()
c. str
d. str[]
Answer #1


ANSWER:-

1) A) 5*7>3*(5+2) && 3*5-4<6

   3*7 = 35

   3*(5+2) = 21

therefore, 35 > 21 = true

3*5-4 == 11

therefore, 11 < 6 = true

   true && true = true

SO, ANSWER is true.

2) .java

3) Point is not a primitive data type. It is a class

4) ANSWER is 7. Because vari[4].length gives no of columns in that row.

5) c. import the package of which the class you are using is a part

6) !( x / 2 == 12 ) || ( z*3 == 96 );

   Explanation:- x = 25

       x/2 = 25/2 = 12

     (x/2 == 12) = true

    !(x/2 == 12) = false

z = 32

   (96 == 96) = true

    !( x / 2 == 12 ) || ( z*3 == 96 ) = false || true = true

ANSWER is true.

7) ANSWER is !. because string length is 15. and indices are in range 0 to 14. and 14th index is !.

8) 5, 3

9) b) the keyword static

10) c. arr[arr.length – 1]

11) arr[9]

12) rand.nextInt(4)+2; rand.nextInt(n) generates random numbers in range 0 to n-1

13) their scopes do not overlap

14) The memory location where array arr is stored.

15) str.length()

NOTE:- If you have any doubts,please comment below.Please give positive rating.THUMBS UP.

          THANK YOU!!!!